OpenStudy (anonymous):

The cost of producing x hundred units of a product is C(x)=3x+5. The revenue earned by producing and selling x units of the product is R(x)=-0.25x2+18x, where x is in hundreds of units and C(x) and R(x) are in thousands of dollars. What will be the profit earned if 300 units are produced and sold?

OpenStudy (anonymous):

C(x)=3(300)+5 R(x)=-.025(300^2)+18(300) R(x)-C(x)=18005